The gross heating value (GHV) of gaseous propane at 298 K is 2219.71 kJ/mol. Calculate the net heating value (NHV) of the propane. Given that, the latent heat of water vapours at 298 K=2442.5 kJ/kg.
1
Expert's answer
2020-12-09T04:59:04-0500

Basis : 1 mol of gaseous propane.


"C_3H_8+5O_2\to 3CO_2 +4H_2O"


"1 mol\space C_3H_8 =4\space mole H_2O"


"i.e\space when \space 1mol \space of \space C_3H_8 \space burnt , 4\space moles \space of \space H_2O \space are \space produced."


Water produced = 4*18 = 72 g = "72 *10^{-3} g"


"NHV = Net \space heating \space value= GHV- Latent \space heat \space of water \space vapor"


"\space\space\space\space\space\space\space\space\space\space\space=2219.71-72*10^{-3}*2442.5"


"\space\space\space\space\space\space\space\space\space\space\space=2043.85\space kJ/mol"